# Webly-Supervised Image Manipulation Localization via Category-Aware Auto-Annotation

### v1

@@ Abstract
Index Terms:
Image manipulation localization, automatic annotation, web supervision, image forensics.
@@ I Introduction
The misuse of manipulated images can lead to fraud and the spread of misinformation, posing serious threats to social media security [1]. To this end, Image Manipulation Localization (IML) has gained increasing attention in recent years as a critical research problem [2, 3]. A central challenge in IML lies in the scarcity of large-scale high-quality data [4], which is essential for preventing model overfitting. Generating such data is extremely labor-intensive and time-consuming [5, 6]. It requires not only crafting convincing manipulated images but also providing precise pixel-level annotations of the forged regions. Synthetic data has been explored as an alternative [7, 8], but the resulting images often differ substantially from real-world forgeries, creating a significant domain gap that hinders model generalization [4]. Recent methods have attempted to close this gap by leveraging noise domain features [9, 10] and semantic suppression techniques [6], showing progress in improving robustness. Nevertheless, these methods do not fully resolve the fundamental issue: the lack of large-scale, manually forged data. As a result, overfitting remains a persistent challenge in IML.
Figure 1: When an image is tampered with by copy-pasting foreground objects from original images, the resulting original-tampered image pair belongs to Shared Donor Group (SDG). When an image is forged by directly modifying elements in place, the pair belongs to Shared Probe Group (SPG).
Figure 2: We propose a novel CIML method that processes the SPG and SDG images separately. It is used for automatic annotation, constructing a large-scale, high-quality dataset. The dataset notably enhances the generalization of image manipulation localization models. Additionally, the Object Jitter method is introduced to further enlarge the dataset size and data diversity.
@@
Meanwhile, there are abundant manually forged images publicly available on the web, often accompanied by their corresponding originals. We propose to leverage such resources to address the data scarcity problem in IML. Specifically, we propose leveraging Constrained Image Manipulation Localization (CIML) models to automatically generate pixel-level mask annotations for these unlabeled forged images. Our key insight is that CIML simplifies the manipulation localization task by comparing a forged image with its corresponding original. This reduction in task complexity allows CIML to produce more accurate and reliable annotations for real-world manipulations. As a result, CIML offers a practical solution for automatically generating large-scale, high-quality training datasets for IML.
However, despite recent progress on less challenging data, existing CIML methods remain inadequate as reliable automatic annotators for complex real-world images, due to three major limitations. First, most prior approaches rely on a single correlation-based model to handle all input data [11, 12], which we argue is a suboptimal design. In practice, image pairs consisting of forged images and their corresponding original images can be categorized into two distinct types based on their relationship (see Fig. 1): (1) Shared Donor Group (SDG), where forged image is created by copying objects from the original image and pasting them onto another image. In SDG, the shared content between the real and fake images is the foreground object. (2) Shared Probe Group (SPG), where forged image is created by directly modifying the original image. In SPG, the shared content is the less object-specific background. Although correlation-based methods work reasonably well on SDG data, they are less effective on SPG data. This is because the shared backgrounds in SPG cover larger areas and offer fewer discriminative features compared to the localized, distinctive foregrounds in SDG. Training a single correlation model on both groups can therefore lead to confusion, resulting in degraded generalization and performance. Second, existing CIML methods overlook an important signal, that is the difference maps obtained by subtracting forged images from their originals. These difference maps consistently highlight manipulated regions and offer strong localization cues on SPG data. However, previous studies did not treat SPG and SDG images separately, thereby leaving this valuable signal underutilized. Third, CIML on SDG is inherently much more challenging than on SPG. Unlike SPG, CIML on SDG cannot benefit from image difference information. Additionally, SDG exhibits greater visual complexity and diverse manipulation patterns. Unfortunately, high-quality training data for SDG is extremely limited, making it difficult to train robust and generalizable models for real-world scenarios.
To tackle these problems, we propose a novel paradigm termed as Category-Aware Auto-Annotation v2 (CAAAv2), which treats image pairs in SDG and SPG separately. Our CAAAv2 comprises three components. At the beginning, a classifier is employed to determine whether an input image pair belongs to SDG or SPG. This classifier can be trained effectively through self-supervised learning using unlabeled images. Moving forward, a Difference-Aware Semantic Segmentation model that utilizes both the image pairs and their difference maps for accurate constrained manipulation localization in SPG. Additionally, a Correlation DINO model that improves generalization on SDG with a frozen DINOv2 [13] encoder, a Learnable Aggregation module, a Feature Super Resolution module and a Multi-Aspect Denoiser.
The proposed CAAAv2 is an upgraded version of our CAAA at CVPR 2024 [14]. In our earlier work, overfitting occurred in SDG because a trainable backbone was used during training with limited data. To address this issue, we now employ a frozen ViT backbone, which has resulted in significant improvements for CIML on SDG.
@@
Subsequently, we collect a large amount of manually forged images from the Internet and then annotate their forged regions with the proposed CAAAv2. This approach considerably alleviates the data scarcity in image manipulation localization. To ensure the reliability of the annotations, we propose Quality Evaluation Score (QES), a novel metric that automatically evaluates annotation quality and filters out inadequate ones. Experiments demonstrate that our MIMLv2 dataset significantly improves generalization of various models on widely used benchmarks, as shown in Fig. 2.
Furthermore, the enhanced precision of CAAAv2 ensures that a greater number of high-quality samples are retained after filtering. Consequently, MIMLv2 is both larger and more diverse than our conference version MIMLv1 [14], better supporting the development of IML models.
To further enhance the scale and diversity of web supervision, we propose Object Jitter. This method generates semantically reasonable high-quality artifacts by introducing slight distortions to random objects, such as enlargement, overexposure, or texture alteration. Then, we train IML models to localize these distorted objects. Our approach effectively supplements MIMLv2 with additional authentic web images. This also differs from our conference paper that only uses MIMLv1 dataset to improve IML models.
@@
To best leverage our high-quality web-scale supervision for IML, we further propose a new model named Web-IML. It comprises an encoder, a Multi-Scale Perception module, and a Self-Rectification module. The Multi-Scale Perception module enables the model to better integrate information from multiple perspectives. The Self-Rectification module allows the model to learn to identify and correct its mistakes. Our Web-IML significantly outperforms previous methods on various widely used benchmarks, being smaller yet more effective than the APSC-Net in our conference version.
@@
  • We propose to address the challenging IML task through the abundant web forgeries and their high-quality annotations distilled from a less challenging task, CIML.
  • We introduce a novel method CAAAv2 for CIML, along with a novel metric QES designed to assess prediction quality. Together, they are used to automatically annotate manually forged web images, constructing a large-scale, diverse, high-quality dataset MIMLv2. This dataset considerably alleviates the data scarcity in IML.
  • To further increase the size and diversity of the training data, we introduce Object Jitter, which generates high-quality artifacts using authentic web images.
  • We design a new model Web-IML, to better leverage the web-scale supervision. Extensive experiments have confirmed the effectiveness of the proposed method.
Figure 3: The overall pipeline of our proposed CAAAv2 paradigm. The input original-manipulated image pairs are first fed into a classifier to be classified into SPG or SDG. Then, we use Difference Aware Semantic Segmentation and Correlation DINO to predict forged regions in SPG and SDG respectively.
@@ II Related Works
Image Manipulation Localization aims to identify forged regions within images [15, 16, 17, 18, 19] . Due to costly data collection, existing handcrafted datasets are tiny in size, leading to overfitting in many models [6, 4]. To address overfitting, some studies incorporated handcrafted features. RGB-N [9] utilized SRM [9] filters to suppress semantic bias. MVSS-Net [20] and MGQFormer [21] further exploited learnable noise domain modeling. CAT-Net [7], ObjectFormer [22] and HiFi-Net [23] leveraged frequency features to localize manipulations. However, the handcrafted features are mostly noisy and unstable, hindering further improvements in model performance. Recently, pure vision models with semantic suppression techniques have achieved better performance [24, 25]. NCL [4] employed contrastive learning to improve generalization. SparseViT [6] utilized dilated attention to reduce semantic bias. Despite the progress made, these works still suffer a lot from over-fitting due to scarce high-quality data. To this end, we propose to address data scarcity by constructing a large-scale high-quality dataset, with the Internet images accurately annotated in an automatic way.
Constrained Image Manipulation Localization aims at identifying the forged image region with reference to its original image [26]. Most prior works relied on correlation matching and treated SDG and SPG image pairs uniformly. DMVN [26] computed correlation maps to localize similar objects across image pairs. DMAC [11] incorporated dilated convolution for enhanced spatial information. AttentionDM [12] employed attention mechanism for better performance. MSTAF [27] performed correlation in both the encoder and decoder to extract better features. While these methods achieved notable progress on less challenging datasets (e.g., synthetic COCO [26]), their performance is limited in real-world scenarios characterized by high resolution, large variations, and complex details. In contrast to previous approaches, our method treats SPG and SDG separately, leveraging the differences between forged images and their originals. We are also the first to employ a frozen backbone for this task, effectively reducing overfitting.
@@ III Category-Aware Auto-Annotation v2
For constrained image manipulation localization (CIML), previous works didn’t consider the critical discrepancy between SPG and SDG images, and processed them using a single correlation-based model. We argue that such a paradigm is sub-optimal and the reasons are as follows:
First, the similar regions in SDG images are typically foreground objects (e.g., the cats in the SDG branch of Fig. 3) that have specific, similar shapes and distinctive features. In contrast, the similar regions in SPG images are background, which generally lacks distinctive features for accurate correlation matching (e.g., for the images of the SPG branch in Fig. 3, a patch of snow in the background is similar to all other snow patches). These background regions will cause confusion for correlation-based models, especially in complex scenarios.
Second, the difference between the forged image and its original in an SPG pair serves as an important cue. Most areas in an SPG image pair are almost identical and spatially aligned (e.g., the image pair in the SPG branch of Fig. 3). Simply subtracting between them yields a difference map that can always highlight the manipulated regions. However, this information is difficult to utilize in correlation-matching-based models and was therefore overlooked in prior works.
Based on these observations, we propose a novel paradigm for the CIML task, Category-Aware Auto-Annotation v2 (CAAAv2). The key idea is to process SPG and SDG images independently, as shown in Fig. 3. First, a classifier (described in Section III.A) categorizes the input image pairs as either SPG or SDG. For SPG pairs, we apply Difference-Aware Semantic Segmentation (Section III.B). For SDG pairs, we process the images using Correlation DINO (Section III.C). Importantly, the trained models are further used to auto-annotate unlabeled forged images, thereby addressing the data scarcity issue in IML tasks.
The difference between CAAAv2 and its conference version CAAA lies in the SDG branch model. In CAAA, we processed SDG image pairs using a trainable CNN structure, which tended to overfit the limited training data. In contrast, CAAAv2 employs a well pre-trained frozen ViT backbone, thereby reducing overfitting to task-specific data. Moreover, the ViT backbone naturally integrates global information, thereby alleviating semantic misalignment caused by scale diversity.
III-A Self-Supervised Classification
@@
To achieve the classification of SPG and SDG, we propose training a classifier via self-supervised learning on unlabeled images. Given an image, we apply random augmentations (e.g. color jitter, image compression) and manipulations (copy-paste, splicing or removal). The manipulated image and the original image form an SPG pair. To construct an SDG pair, we copy 1-3 random regions from the original image, resize them, and paste them into another image. With the obtained image pairs, we can effectively train our classifier. Our classifier actually only needs to discern whether the main area of the input image pair is identical or not, without considering which image or region is fake. For example, in Fig. 8, the two images in any SPG pair are nearly identical and spatially aligned, whereas images in any SDG pair are distinctly different. These characteristics are universally applicable and strong enough. Therefore, this classification task is quite simple, enabling accurate separation of images into the two groups.

### v2

@@ Abstract
Figure 1: Left: Definition and characteristics of the Shared Donor Group (SDG) and Shared Probe Group (SPG). Right: In contrast to previous works, our CAAAv2 adopts category-aware and prior-feature-denoising paradigm that notably reduces task difficulty and overfitting. We also propose using the CAAAv2 models to auto-annotate web-scale manually forged images. This approach significantly addresses the data scarcity issue for IML and notably improves the generalization of all IML models.
footnotetext: This research is supported in part by NSFC (Grant No.: 62476093) and GD-NSF (No.2021A1515011870). (Corresponding author: Lianwen Jin.) Chenfan Qu, Huiguo He, and Lianwen Jin are with the School of Electronics and Information Engineering, South China University of Technology, Guangdong 510641, China. (Email: 202221012612@mail.scut.edu.cn, hehuiguo@scut.edu.cn, eelwjin@scut.edu.cn) Yiwu Zhong is with the School of Intelligence Science and Technology, Peking University, Beijing 100091, China. (Email: yiwu-zhong@outlook.com) Bin Li is with the College of Electronics and Information Engineering, Shenzhen University, Shenzhen 518060, China. (Email: libin@szu.edu.cn)
@@ I Introduction
The misuse of manipulated images can lead to fraud and the spread of misinformation, posing serious threats to social media security [1, 2, 3, 4, 5, 6, 7, 8]. To this end, Image Manipulation Localization (IML) has garnered increasing attention in recent years [9, 10, 11, 12, 13, 14, 15, 16]. A central challenge in IML lies in the scarcity of large-scale high-quality data [17], which is essential for preventing model overfitting. Creating such data is extremely labor-intensive and time-consuming [18], requiring not only the crafting of convincing forgeries but also the precise pixel-level annotation of manipulated regions. Although alternatives like synthetic data [19, 20] or AI-generated forgeries [21] have been explored, they often exhibit significant discrepancies from real-world manipulations, thereby limiting model generalization to practical scenarios [17].
@@
Meanwhile, numerous unlabeled, manually forged images are publicly available over the Internet, often accompanied by their corresponding originals. To address the scarcity of annotated data, we propose to leverage these unlabeled forgeries by automatically generating pixel-level mask annotations using Constrained Image Manipulation Localization (CIML) models. Our key insight is that CIML can reduce the complexity of manipulation localization by comparing the forged image with its original, resulting in more accurate and reliable annotations. Therefore, CIML offers a practical solution for automatically generating large-scale, high-quality training datasets for IML.
However, despite recent progress with CIML on simple synthetic images, existing methods are unreliable for complex, real-world forgeries due to two major limitations:
First, previous methods employ a suboptimal paradigm that relies on a single correlation-based model for all input types [22, 23]. In practice, image pairs consisting of forged images and their original counterparts can be categorized into two distinct groups (see Fig. 1): (1) Shared Donor Group (SDG) denotes image pairs that share foreground, with forged images generated by copying objects from the original image; (2) Shared Probe Group (SPG) denotes image pairs that share background, where forged images are obtained by modifying the original image in-place. While correlation-based methods perform reasonably well on SDG data, they falter on SPG data because the shared region (the background) often lacks distinctive features. Forcing a single model to handle both groups leads to confusion and degraded performance.
Second, the CIML task itself is challenging and also suffers from a scarcity of high-quality training data. Previous works provide the model with only the image pair, relying entirely on the network to learn the complex correspondence between the images in each pair. Without clear and explicit guidance, models struggle to learn this correspondence effectively and are prone to overfitting on limited training data.
To overcome these challenges, we propose a novel paradigm: Category-Aware Auto-Annotation v2 (CAAAv2). It achieves superior generalization via two key innovations:
(1) Category awareness. We treat image pairs in SDG and SPG separately through a classifier at start. This separation avoids the confusion of joint training and allows us to deploy distinct, optimized pipelines that leverage the unique correspondence of each category. This classifier can be effectively trained on unlabeled images.
(2) Task transformation. We reframe the challenging CIML problem as a far simpler denoising task. This is achieved by introducing novel prior features that provide strong initial guidance for localization. Crucially, these prior features are extracted via zero-shot frozen functions that are resistant to overfitting. For SDG pairs, where the forged regions are typically shared objects, we extract features from both images using a frozen DINOv2 encoder and compute their correlation map. The resulting map, named DINO correlation, serves as a prior that roughly highlights the manipulated regions (Fig. 1). The model’s task is reduced to simply denoising this prior to produce a precise mask. For SPG pairs, where the forged regions are the difference between the images, we obtain the prior map by subtracting the forged image from its original. Through task transformation, we significantly reduce task difficulty and bridge the gap between synthetic and real-world data, leading to substantially better generalization.
@@
Subsequently, we collect a large amount of manually forged images from the Internet and then annotate their forged regions with the proposed CAAAv2. This approach considerably alleviates the data scarcity in image manipulation localization. To ensure the reliability of the annotations, we propose Quality Evaluation Score (QES), a novel metric that automatically assesses and filters out inadequate annotations.
While this pipeline yields a high-quality dataset, we observed a relative lack of sophisticated copy-move forgeries that feature only subtle edge artifacts. Direct copy-move synthesis is not effective since it brings obvious semantic artifacts. To address this, we introduce Object Jitter, a data synthesis method that simulates these challenging cases by applying minor perturbations (e.g., enlargement) to random objects on authentic images, effectively supplementing our dataset.
In our CVPR 2024 paper, we propose to address the data scarcity of IML by auto-annotating images from the web with a category-aware approach CAAA. However, CAAA did not incorporate a prior-feature-denoising paradigm for its SDG branch, leading to overfitting on limited training data. In this paper, we propose CAAAv2, an improved version of our CAAA that resolves this issue by leveraging frozen DINO-based prior features. The enhanced accuracy of CAAAv2 ensures that a greater number of samples are retained after filtering. Consequently, the collected dataset MIMLv2 is both larger and more diverse than our conference version MIMLv1 [24], better supporting the development of IML models (Figure 1). Moreover, different from our conference paper that only uses MIMLv1 dataset to improve IML models, our Object Jitter approach further effectively supplements MIMLv2 with high-quality hard samples.
@@
To better leverage our high-quality web-scale supervision for IML, we further propose a new model named Web-IML, which comprises a Multi-Scale Perception module and a Self-Rectification module. The former enables the model to better integrate information from multiple perspectives; the latter allows the model to learn to identify and correct its mistakes. Our Web-IML significantly outperforms previous methods on various widely used benchmarks, being smaller yet more effective than the APSC-Net in our conference version.
@@
  • We propose to address the scarcity of IML training data by leveraging the less challenging CIML task to auto-annotate forged images collected from the web.
  • We introduce CAAAv2, a novel CIML paradigm that significantly enhances generalization. This is achieved through a category-aware design, which treats distinct image groups separately, and a prior-denoising approach, which reframes CIML as a simpler denoising task.
  • With CAAAv2 and a novel metric QES, we automatically annotate manually forged web images, constructing a large-scale, diverse, high-quality dataset MIMLv2. This dataset considerably alleviates the data scarcity in IML.
  • We introduce Object Jitter, which generates high-quality, challenging-to-localize artifacts, effectively supplementing our dataset. To better leverage the web-scale supervision, we design a novel model Web-IML. Extensive experiments have validated its effectiveness.
Figure 2: The overall pipeline of our proposed CAAAv2 paradigm. The input original-forged image pairs are first categorized as either SDG or SPG with a classifier. Then, we use Corr-DINO and Difference-Aware Semantic Segmentation to process SDG and SPG pairs respectively. Both models first extract a prior feature that roughly highlights forgeries and then denoise it.
@@ II Related Works
II-A Image Manipulation Localization (IML)
Image Manipulation Localization aims to identify forged regions within images [25, 26, 27, 28, 29, 30, 31, 32, 33]. The scale of existing handcrafted datasets is limited due to the high cost of manual collection and annotation, inevitably leading to overfitting [18, 17]. To mitigate this issue, some studies have incorporated handcrafted features. For instance, RGB-N [34] applied SRM [34] filters to suppress semantic bias. MVSS-Net [35] and MGQFormer [36] further exploited learnable noise domain modeling. CAT-Net [19], ObjectFormer [37] and HiFi-Net [38] leveraged frequency features to localize manipulations. However, the handcrafted features are noisy and unstable, hindering further improvements in model performance. Recently, pure vision models have made notable progress [39, 1]. NCL [17] employed contrastive learning to improve generalization. SparseViT [18] utilized dilated attention to reduce semantic bias. PIM-Net [1] explicitly modeled the super-pixel inconsistency to alleviate biased learning. Despite these advances, overfitting remains a significant challenge due to the fundamental issue: the scarcity of high-quality training data.
To address this fundamental issue, we propose to construct a large-scale, high-quality dataset by automatically annotating Internet images with a less challenging task, CIML.
II-B Constrained Image Manipulation Localization (CIML)
Constrained Image Manipulation Localization aims to identify forged regions in an image by leveraging its original counterpart [40]. Previous works relied on correlation matching and handled SDG and SPG image pairs uniformly. For instance, DMVN [40] computed correlation maps to localize similar objects across image pairs. DMAC [22] incorporated dilated convolution to enhance spatial information. AttentionDM [23] employed an attention mechanism to improve performance. MSTAF [41] performed correlation in both the encoder and decoder to extract richer features. While these methods achieved notable progress on less challenging datasets (e.g., synthetic COCO [40]), their performance is limited in real-world forgeries, which often involve high-resolution images, significant variations, and complex details.
In contrast, our approach reduces task difficulty and addresses overfitting by (1) explicitly harnessing the distinct characteristics of SPG and SDG pairs; and (2) transforming the CIML problem into a denoising task through prior features extracted from frozen functions.
@@ III-A Motivation and Framework Overview
For CIML, previous works overlooked the critical discrepancy and characteristics between SPG and SDG images, treating them using a single fully-tunable correlation model. We argue that this paradigm is suboptimal for two reasons:
1. Conflicting Correlation Signals. The nature of the ”shared region” differs drastically between the two categories. In SDG pairs, the shared region is salient foreground objects with distinctive features (e.g., the lion in SDG of Fig. 2), providing a strong signal for correlation matching. Conversely, in SPG pairs, the shared region is the background, which lacks the unique features required for reliable matching (e.g., the sky in SPG of Fig. 2). Forcing a single correlation model to simultaneously handle both groups creates a conflicting objective that degrades performance.
2. Lack of Explicit Guidance. Learning to comprehend and fully leverage the correspondence between original and manipulated images is non-trivial. With scarce training data, CIML models tend to find shortcuts to minimize training loss, often overfitting to either forgery-agnostic semantics in small-scale handcrafted datasets or monotonous patterns in synthetic data. Without explicit guidance on how to utilize the correspondence within an image pair, models often fail to generalize to diverse, real-world manipulations.
To overcome these challenges, we propose Category-Aware Auto-Annotation v2 (CAAAv2), a novel paradigm for CIML. Our core idea is to process SDG and SPG images independently, while explicitly integrating image correspondence as a guiding prior feature. Specifically, as shown in Fig. 2, CAAAv2 introduces a classifier to categorize the input as either SPG or SDG, and then applies a category-specific prior to guide localization. For SDG pairs, we introduce DINO Correlation, a prior that captures shared foreground objects by correlating dense features from a frozen DINOv2 encoder. For SPG pairs, a simple Difference Map effectively identifies manipulated regions by subtracting the two images.
These priors embed domain-specific knowledge about forgery patterns into strong, zero-shot signals. By generating these priors with frozen, non-trainable functions, we reframe the complex task of open-ended forgery localization into a simpler, more constrained denoising problem. Here, the model’s objective is to refine the coarse prior into a precise mask, using the input images as context. This paradigm shift significantly reduces task difficulty and minimizes the generalization gap between synthetic and real-world forgeries, leading to substantial performance improvements.
Importantly, the trained models are further used to auto-annotate large-scale, unlabeled manual forgeries from the web, thereby addressing the data scarcity issue in IML.
Whereas CAAA overfits on SDG data, CAAAv2 provides explicit guidance by introducing DINO Correlation as a strong prior. This prior is generated using a frozen DINOv2, whose global receptive field naturally aligns multi-scale semantics, yielding substantial gains in real-world generalization.
III-B Self-Supervised Classification
@@
Training a classifier to distinguish between SDG and SPG is a non-trivial task due to the lack of labeled data. To address this challenge, we propose training a classifier via self-supervised learning on unlabeled images. Given an image, we apply random augmentations (e.g., color jitter, image compression) and in-place manipulations (copy-paste, splicing, or removal). The manipulated image and the original image form an SPG pair. To construct an SDG pair, we copy 1-3 random regions from the original image, resize them, and paste them into another image. With the obtained image pairs, we can effectively train our classifier. Our classifier actually only needs to discern whether the main area of the input image pair is identical or not, without considering which image or region is fake. For example, in Fig. 6, the two images in any SPG pair are nearly identical and spatially aligned, whereas images in any SDG pair are distinctly different. These characteristics are universally applicable and strong enough. Therefore, this classification task is quite simple, enabling accurate separation of images into the two groups.
III-C Corr-DINO
For Shared Donor Group (SDG) pairs, the manipulated regions are mostly the objects copied from the original image. The localization task therefore reduces to identifying these shared foreground objects. To exploit this correspondence, we introduce a novel prior feature, termed DINO Correlation. This prior is generated by extracting dense features from both images using a frozen DINOv2 encoder and then computing their cross-correlation. The resulting map provides a strong yet coarse signal that highlights the shared objects, which often are the forged regions. The task is thus simplified into refining this coarse prior to a precise segmentation mask. Based on this, we propose the Corr-DINO model. As illustrated in Figure 3, it consists of four components:
1. Prior Feature Extractor for prior feature extraction. We extract object features from both the tampered image $I_{a}$ and its original image $I_{b}$ with a frozen DINOv2 [42] backbone, which has learned generalized representations through self-supervised learning on vast-scale web images. The features from the last DINO layer, denoted as $F_{a1}$ and $F_{b1}$, contain rich semantic information and are used to compute correlation features $F_{a\_corr}$ and $F_{b\_corr}$. $F_{a\_corr}=[Corr(F_{a1},F_{b1}),Corr(F_{a1},F_{a1})]$, $F_{b\_corr}=[Corr(F_{b1},F_{a1}),Corr(F_{b1},F_{b1})]$. The same correlation function $Corr$ as most previous works [23, 41, 24] is adopted.
Figure 3: The overall pipeline of our proposed Corr-DINO. First, we employ a frozen DINO backbone to extract image features. The features obtained from the final DINO layer are then used for correlation calculation. Next, a learnable aggregation module reduces the channel dimension of the correlation features, which are further refined using a Feature Super Resolution module. Finally, the refined features are denoised with a Multi-Aspect Denoiser, resulting in the final mask prediction. Green lines indicate operations used only during training with synthetic data, due to the lack of mask labels for the real images.
2. Learnable Aggregation module for dimension reduction. Previous works simply select the top-K activated channels with the highest average values in $F_{a\_corr}$ and $F_{b\_corr}$ to reduce dimension. This approach works for trainable encoders but is suboptimal for our frozen encoder, whose correlated object features are relatively sparse. Selecting only the top-K channels may discard important features and compromise the integrity of $F_{a\_corr}$ and $F_{b\_corr}$. To overcome this, we propose a learnable aggregation module that uses trainable parameters to reduce the channel count to K. We also include statistics of average and maximum values across the channel dimension. $F_{a\_aggr}=[C_{1\times 1}(R(C_{1\times 1}(F_{a\_corr}))),Avg(F_{a\_corr}),Max(F_{a\_corr})]$. Here, $C_{1\times 1}$ is $1\times 1$ conv-layer, $R$ is ReLU. $Avg$ and $Max$ are average and maximum pooling respectively. A similar process is used for $F_{b\_corr}$ to generate $F_{b\_aggr}$.
3. Feature Super Resolution module for image detail reconstruction. The features from the last DINOv2 layer are rich in semantics but lack fine image details [42], making $F_{a\_aggr}$ and $F_{b\_aggr}$ too coarse for fine-grained mask predictions. To reconstruct image details, we leverage features from the last 2-4 layers of DINO, which are known to be more sensitive to edges and fine details [42]. We concatenate the features from the last four encoder layers $F_{a}=Concat(F_{a1,2,3,4})$ and interpolate $F_{a\_aggr}$ to different scales ($1\times,2\times,4\times,8\times$) to obtain $F_{aggr\_a1,2,3,4}$. We then fuse them with $1\times 1$ conv-layers. A similar process is applied to $F_{b\_aggr}$.
4. Multi-Aspect Denoiser for feature denoising. Since the DINOv2 encoder is frozen, its output features inevitably contain significant task-irrelevant noise. To achieve precise mask prediction, we propose denoising the correlation features by integrating multi-perspective information. This module fuses multi-scale features $F_{asr\_1,2,3,4}$ in a top-down manner. The fused features are further concatenated and denoised using a set of dilated conv-layers. The output is concatenated to $F^{\prime}1$, which is used to yield the mask prediction. More details are provided in the Appendix.